# Graph Signal Processing for Global Stock Market Volatility Forecasting

### v1

@@ 4.2 Magnetic Laplacian
A complex Hermitian matrix $\mathbf{H}$ is a square matrix such that it is equal to its conjugate transpose: $\mathbf{H}=\mathbf{H}^{\dagger}$, which means $\mathbf{H}_{ij}=\overline{\mathbf{H}_{ji}}$ and $\overline{\mathbf{H}_{ji}}$ is the complex conjugate of element $\mathbf{H}_{ij}$. Furthermore, the unnormalized and normalized magnet Laplacians can be formulated in Equations (17) and (18), respectively,
@@
Different $q$ values represent different handlings of the directionality of the graph. For example, if $q=0$, $\mathbf{H}^{(0)}=\mathbf{A}^{s}$. The directionality of the graph is eliminated through symmetrizing the asymmetric adjacency matrix. If $q=0.25$, $\mathbf{L}_{m}^{(0.25)}=-(\mathbf{L}_{m}^{(0.25)})^{\top}$ when there is only one directed edge between any two connected nodes of $\mathcal{G}$.
Take the normalized magnet Laplacian matrix $\widetilde{\mathbf{L}}_{m}^{(q)}$ as an example. Its eigendecomposition can be formulated in Equation (19) below:
@@
where columns in $\mathbf{U}_{m}$ are still the orthonormal basis of eigenvectors of $\widetilde{\mathbf{L}}_{m}^{(q)}$ and $\mathbf{U}_{m}^{\dagger}$ is the conjugate transpose of $\mathbf{U}_{m}$. It has been proved that $\mathbf{U}_{m}$ and $\mathbf{U}_{m}^{\dagger}$ are complex valued while eigenvalues in $\bm{\Lambda}_{m}$ are all real non-negative numbers (Zhang et al., 2021). This allows easy extensions of classical GSP techniques to directed graphs. The GFT and IGFT for graph signal $\mathbf{x}\in\mathbb{R}^{N}$ based on normalized magnet Laplacian are formulated in Equations (20) and (21), respectively,
@@
The magnet Laplacian offers a powerful tool to extend the GSP techniques to directed graphs. By adjusting the value of $q$, the magnet Laplacian can suit both the undirected and direct graph structures and can process the directional information in the graphs as needed. When $q>0$, the eigenvalue decomposition of the magnet Laplacian leverages the graph directional information by extending the graph Fourier bases to the complex domain, while the eigenvalues remain non-negative real numbers as meaningful graph signal frequency measurements. This is extremely useful in multivariate RV forecasting problems in that the volatility spillover effect between different markets can have different scales and directions (Diebold and Yilmaz, 2012). This means that, compared to an undirected or unweighted graph, a directed and weighted graph is more suitable to capture the volatility spillover network consisting of various global stock markets. In this case, compared to the GNNHAR model relying on spatial GNNs which are biased in processing directional information, models applying GSP techniques based on the magnet Laplacian are more powerful in learning the volatility dynamics (Zhang et al., 2021). The details on the construction of the directed and weighted volatility spillover graph are presented in the next section.
4.3 Volatility Spillover Graph Formulation
There are various ways to formulate the volatility spillover graph for global stock market indices based on their historical RV observations. For example, the Pearson correlation method and the Graphical LASSO (GLASSO) method (Friedman et al., 2007) can produce a symmetric matrix to represent the connections between the RV time series of different stock market indices. However, these methods cannot capture the imbalanced impact between influential stock markets and impressionable stock markets. On the other hand, the Diebold and Yilmaz (DY) framework can capture volatility spillovers from and to specific markets (Diebold and Yilmaz, 2012).
Suppose a covariance stationary multivariate RV time series $\{\mathbf{v}_{t}\in\mathbb{R}^{N}\}_{t=1}^{T}$ consists of $N$ different stock market indices. A $p$-order vector autoregressive model (VAR($p$)) for these time series can be represented as:
$\mathbf{v}_{t}=\sum_{i=1}^{\infty}\mathbf{B}_{i}\bm{\epsilon}_{t-i},$ (23)
For this study, the DY framework is preferred because it incorporates directionality. Here, the hyperparameters $p=22$ to align the HAR model and $H$ is set according to specific forecast tasks. Besides, the net pairwise volatility spillover effect matrix $\overline{\bm{\theta}}^{g}(H)$ is applied because it can effectively highlight the dominant channels of volatility transmission, which allows researchers to pinpoint the markets that exert the most significant influence over others.
@@ 4.4 Model Formulation
In this section, the proposed vanilla GSPHAR (v-GSPHAR for short) and GSPHAR models are introduced. We first present the simpler and more basic v-GSPHAR model to demonstrate how GSP techniques are integrated into the HAR framework for RV forecasting. After that, a more flexible and sophisticated GSPHAR model is developed. The general workflow of the GSPHAR model is described in Figure 2 below:
@@
Now we present the detailed step-by-step development of the proposed framework. Suppose that the multivariate RV time series $\{\mathbf{v}_{t}\in\mathbb{R}^{N}\}_{t=1}^{T}$ consist of the RV time series of $N$ different stock market indices from time $t=1$ to time $t=T$. In the v-GSPHAR model, we eliminate the directional information in the volatility spillover graph by setting the hyperparameter $q=0$ in the magnet Laplacian (18). Thus, the Hermitian adjacency matrix $\mathbf{H}^{(q)}$ and the symmetric adjacency matrix $\mathbf{A}^{s}$ become identical and they are constructed based on the net pairwise volatility spillovers $\overline{\bm{\theta}}^{g}(H)$:
$\mathbf{A}^{s}=\frac{1}{2}(\mathbf{A}+\mathbf{A}^{\top})=\frac{1}{2}(\overline{\bm{\theta}}^{g}(H)+\overline{\bm{\theta}}^{g}(H)^{\top}).$ (27)
The normalized magnet Laplacian matrix for the v-GSPHAR model is formulated as:
$\widetilde{\mathbf{L}}_{m}^{(0)}=\mathbf{I}-{(\mathbf{D}^{s})}^{-\frac{1}{2}}\mathbf{A}^{s}{(\mathbf{D}^{s})}^{-\frac{1}{2}}.$ (28)
$\widetilde{\mathbf{L}}_{m}^{(0)}$ is symmetric and positive semidefinite. Hence, it can be decomposed into real matrices through eigendecomposition as in Equation (10) and the multivariate RV time series can be projected to the spectral domain through the GFT as shown below:
$[\widetilde{\mathbf{v}}_{t-1},\widetilde{\mathbf{v}}_{t-2},\ldots,\widetilde{\mathbf{v}}_{t-22}]=\mathbf{U}_{m}^{\top}[\mathbf{v}_{t-1},\mathbf{v}_{t-2},\ldots,\mathbf{v}_{t-22}]$ (29)
where $\widetilde{\mathbf{L}}_{m}^{(0)}=\mathbf{U}_{m}\bm{\Lambda}_{m}\mathbf{U}_{m}^{\top}$. Hence, a set of basis-dependent univariate HAR models can be constructed in the graph Fourier domain:
${\widetilde{\mathbf{v}}_{t}}=\bm{\alpha}+\bm{\beta}_{d}\widetilde{\mathbf{v}}_{t-1}+\bm{\beta}_{w}\widetilde{\mathbf{v}}_{t-5:t-1}+\bm{\beta}_{m}\widetilde{\mathbf{v}}_{t-22:t-1},$ (30)
where $\widetilde{\mathbf{v}}_{t-n:t-1}=\frac{1}{n}\sum_{j=t-n}^{t-1}\widetilde{\mathbf{v}}_{j}$ is the average spectral RV data of each graph Fourier basis from time $t-1$ to time $t-n$. In addition, $\bm{\alpha}$, $\bm{\beta}_{d}$, $\bm{\beta}_{w}$ and $\bm{\beta}_{m}$ are all diagonal square matrices containing the graph Fourier basis–specific parameters. The forecast spectral RV values are projected back to the spatial domain to derive the final forecast output $\widehat{\mathbf{v}}_{t}$ through IGFT:
$\widehat{\mathbf{v}}_{t}=\mathbf{U}_{m}{\widetilde{\mathbf{v}}_{t}}.$ (31)
In this way, the model can be trained and the parameters $\bm{\alpha}$, $\bm{\beta}_{d}$, $\bm{\beta}_{w}$ and $\bm{\beta}_{m}$ can be estimated through the loss measured by any loss function with respect to the true RV value $\mathbf{v}_{t}$ and the forecast RV value $\widehat{\mathbf{v}}_{t}$.
In summary, the v-GSPHAR model constructs and trains a set of univariate HAR models in the graph Fourier domain. Compared to the VHAR and HAR-KS models, the number of parameters in the v-GSPHAR model increases linearly with the addition of more stock market indices. Furthermore, the volatility spillover effect is captured through GFT and its inverse (IGFT). This enables the model to perceive a broader impact range compared to methods that rely on spatial neighborhood information passing. However, this v-GSPHAR still holds the linearity assumption in model construction and cannot effectively capture the directional information in the volatility spillover graph. To address these limitations, we propose an improved GSPHAR model below.
@@
Essentially, the improved GSPHAR model can be viewed as a combination of two v-GSPHAR models with learnable average pooling weights, integrated through a shallow neural network. To learn from the directional volatility spillover effect, the hyperparameter $q$ of the magnet Laplacian should be nonzero. Since the net pairwise volatility spillovers $\overline{\bm{\theta}}^{g}(H)$ is applied here, $q=0.25$ is chosen for better clarity and interpretability as $\mathbf{L}_{m}^{(0.25)}=-(\mathbf{L}_{m}^{(0.25)})^{\top}$ when only one directed edge exists between any two connected nodes of the volatility spillover graph. Since $q=0.25$, the eigendecomposition of the magnet Laplacian $\widetilde{\mathbf{L}}_{m}^{(0.25)}$ returns two eigenvector matrices, $\mathbf{U}_{m}$ and $\mathbf{U}_{m}^{\dagger}$, in the complex space and one eigenvalue matrix, $\bm{\Lambda}_{m}$, containing non-negative real numbers in its diagonal as shown in Equation (19). Thus, the multivariate RV time series input $[\mathbf{v}_{t-1},\mathbf{v}_{t-2},\ldots,\mathbf{v}_{t-22}]\in\mathbb{R}^{N\times 22}$ can be projected to the spectral domain through the GFT:
@@
Here, each projected value in $\{\widetilde{\mathbf{v}}_{t-p}\}_{p=1}^{22}$ is complex valued and can be written as: $\widetilde{\mathbf{v}}_{t-p}=\widetilde{\mathbf{v}}^{\text{r}}_{t-p}+i\widetilde{\mathbf{v}}^{\text{i}}_{t-p}$. The projected RV data merges the volatility information between different stock markets under the guidance of the volatility spillover graph. The GSPHAR model is designed to process the information in both the real and imaginary domains in a way similar to the HAR family of models.
The HAR family of models always applies an equal-weight average to construct the volatility pattern features based on the historical RV data. This assumption can be potentially improved. For example, under the simple design of the univariate HAR model in Equation (2), $\text{RV}_{t-2}$ and $\text{RV}_{t-5}$ always exert the same influence on the forecast $\text{RV}_{t}$. This equal weighting may not be able to sufficiently account for the dynamics of financial markets where the influence of past volatility on future volatility typically diminishes over time. This diminishing influence not only exists between short-term, mid-term and long-term, but also possibly occurs within different periods. To handle such potential insufficiency, a set of graph Fourier basis-specific convolution filters with learnable convex weights is applied to aggregate past information in the mid-term and long-term:
@@
Hereafter, two distinct univariate HAR models are built for the real and imaginary domains, respectively. For clarity, let $\widetilde{\mathbf{V}}_{t-22:t-1}=[\mathbf{1},\widetilde{\mathbf{v}}_{t-1},\widetilde{\mathbf{v}}_{t-5:t-1},\widetilde{\mathbf{v}}_{t-22:t-1}]\in\mathbb{R}^{N\times 4}$ to denote a collection of intercept terms and past short-term, mid-term and long-term volatility patterns. The matrix $\mathbf{W}\in\mathbb{R}^{4\times 1}$ contains the linear transformation parameters used in the HAR model. Here, superscript ‘r’ is used to denote the real component and superscript ‘i’ represents the imaginary component. The same set of HAR coefficients is applied to different graph Fourier bases because the different lagged impact is believed to be well captured in the convolution filters defined in Equation (33). For each graph Fourier basis, its RV forecast in the real and imaginary domains can be formulated as:
$\begin{split}\widetilde{\mathbf{v}}^{\text{r}}_{t}&\;=\widetilde{\mathbf{V}}_{t-22:t-1}^{\text{r}}\mathbf{W}^{\text{r}},\\ \widetilde{\mathbf{v}}^{\text{i}}_{t}&\;=\widetilde{\mathbf{V}}_{t-22:t-1}^{\text{i}}\mathbf{W}^{\text{i}}.\end{split}$ (34)
Set $\widetilde{\mathbf{v}}_{t}=\widetilde{\mathbf{v}}^{\text{r}}_{t}+i\widetilde{\mathbf{v}}^{\text{i}}_{t}$, which is still a complex number. $\widetilde{\mathbf{v}}_{t}$ is projected back to the spatial domain through the IGFT:
@@
The complex valued $\overline{\mathbf{v}}_{t}$ is divided into its real and imaginary parts again: $\overline{\mathbf{v}}_{t}=\overline{\mathbf{v}}^{\text{r}}_{t}+i\overline{\mathbf{v}}^{\text{i}}_{t}$. The information of the real part and the imaginary part of $\overline{\mathbf{v}}_{t}$ is aggregated by a $3$ layer neural network $NN(\cdot)$ to output the final forecast: $\widehat{\mathbf{v}}_{t}=NN(\overline{\mathbf{v}}^{\text{r}}_{t},\overline{\mathbf{v}}^{\text{i}}_{t})$. The shallow neural network can merge the volatility information in the real and imaginary domains in a nonlinear way, which effectively enhances the expressiveness of the proposed GSPHAR model. Compared to the v-GSPHAR model, the improved GSPHAR model is more flexible in capturing the directional spillover effect and forecasting future volatility patterns at the cost of increasing parameters. Despite the complex model design, the growth rate of the parameters still remains linear with respect to the number of stock market indices included. This ensures that the GSPHAR model is not only powerful, but also scalable.

### v2

@@ 2.2 GSP Based on the Magnetic Laplacian
A complex Hermitian matrix $\mathbf{H}$ is a square matrix such that it is equal to its conjugate transpose: $\mathbf{H}=\mathbf{H}^{\dagger}$, which means $\mathbf{H}_{ij}=\overline{\mathbf{H}_{ji}}$ and $\overline{\mathbf{H}_{ji}}$ is the complex conjugate of element $\mathbf{H}_{ij}$. Furthermore, the unnormalized and normalized magnetic Laplacians can be formulated in Equations (6) and (7), respectively,
@@
Different $q$ values represent different handlings of the directionality of the graph. For example, if $q=0$, $\mathbf{H}^{(0)}=\mathbf{A}^{s}$. The directionality of the graph is eliminated through symmetrizing the asymmetric adjacency matrix. Thus, if the graph $\mathcal{G}$ is undirected, $q$ is fixed to 00. If $q=0.25$, $\mathbf{L}_{m}^{(0.25)}=-(\mathbf{L}_{m}^{(0.25)})^{\top}$ when there is only one directed edge between any two connected nodes of $\mathcal{G}$.
Take the normalized magnetic Laplacian matrix $\widetilde{\mathbf{L}}_{m}^{(q)}$ as an example. Its eigendecomposition can be formulated in Equation (8) below:
@@
where columns in $\mathbf{U}_{m}$ are the orthonormal basis of eigenvectors of $\widetilde{\mathbf{L}}_{m}^{(q)}$ and $\mathbf{U}_{m}^{\dagger}$ is the conjugate transpose of $\mathbf{U}_{m}$. It has been proved that $\mathbf{U}_{m}$ and $\mathbf{U}_{m}^{\dagger}$ are complex valued while eigenvalues in $\bm{\Lambda}_{m}$ are all real non-negative numbers (Zhang et al., 2021). This allows easy extensions of classical signal processing techniques to graphs.
The Graph Fourier Transformation (GFT) and its inverse (IGFT) for graph signal $\mathbf{x}\in\mathbb{R}^{N}$ based on normalized magnetic Laplacian are formulated in Equations (9) and (10), respectively,
@@
By adjusting the value of $q$, the magnetic Laplacian can suit both the undirected and direct graph structures and can process the directional information in the graphs as needed. When $q>0$, the eigenvalue decomposition of the magnetic Laplacian leverages the graph directional information by extending the graph Fourier bases to the complex domain, while the eigenvalues remain non-negative real numbers as meaningful graph signal frequency measurements.
2.3 HAR Model and Its Extensions on Volatility Spillovers
The HAR framework is designed to account for the fact that different market participants (such as high frequency traders, institutional investors, and long-term investors) operate on different time horizons. The HAR model incorporates pooled RV data from these various time scales, typically daily, weekly and monthly, to represent the volatility behaviors. This design helps to capture the heterogeneous nature of volatility patterns in financial markets. Since the original RV data usually shows leptokurtosis (fat tails) and skewness, the log-transformed and square root versions of the RV data are often utilized to improve the forecasting accuracy of the HAR models.
Suppose $\mathbf{v}_{t}$ contains the square root RV value of each of $N$ stock market indices on day $t$: $\mathbf{v}_{t}=[\left(\text{RV}_{1,t}\right)^{1/2},\left(\text{RV}_{2,t}\right)^{1/2},\ldots,\left(\text{RV}_{N,t}\right)^{1/2}]$ and $\mathbf{v}_{t-n:t-m}$ is the average-pooled data: $\mathbf{v}_{t-n:t-m}=\frac{1}{n-m}\sum_{j=t-n}^{t-m}\mathbf{v}_{j}$. $\widehat{\mathbf{v}}_{t}$ denotes the forecast square root RV value on day $t$. The conventional HAR model is formulated as:
$\widehat{\mathbf{v}}_{t}=\bm{\alpha}+\bm{\beta}_{d}\mathbf{v}_{t-1}+\bm{\beta}_{w}\mathbf{v}_{t-5:t-1}+\bm{\beta}_{m}\mathbf{v}_{t-22:t-1},$ (11)
where coefficients $\bm{\beta}_{d}$, $\bm{\beta}_{w}$ and $\bm{\beta}_{m}$ are diagonal matrices and they measure how the past daily, weekly and monthly volatility patterns influence future RVs. The intercept term $\bm{\alpha}$ is also a diagonal matrix. The HAR model cannot capture the comovement or correlations of the volatility in different stock markets, which limits its ability to deliver more accurate forecasts.
The GNNHAR model adds a $l$-layer spatial GNN component to the HAR model as an additional input variable to capture the volatility spillover effect as the complex interactions between nodes through neighborhood information aggregation and transformation. The GNNHAR model precomputes the volatility relational network $\mathcal{G}$ with an adjacency matrix $\mathbf{A}$ and a degree matrix $\mathbf{D}$ which remain fixed during the training process. The model is formulated as:
$\begin{split}\mathbf{H}^{(0)}=&\;[\mathbf{v}_{t-1},\mathbf{v}_{t-5:t-2},\mathbf{v}_{t-22:t-6}],\\ \mathbf{H}^{(1)}=&\;\text{ReLU}(\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}\mathbf{H}^{(0)}\mathbf{W}^{(0)}),\\ &\ldots,\\ \mathbf{H}^{(L+1)}=&\;\text{ReLU}(\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(L)}),\\ \widehat{\mathbf{v}}_{t}=&\;\bm{\alpha}+\beta_{d}\mathbf{v}_{t-1}+\beta_{w}\mathbf{v}_{t-5:t-2}+\beta_{m}\mathbf{v}_{t-22:t-6}\\ &\;+\bm{\gamma}\mathbf{H}^{(L+1)}+\bm{\epsilon}_{t}.\end{split}$ (12)
Here, $\{\mathbf{W}^{(l)}\}_{l=0}^{L}$ are the trainable graph information transformation parameters, representing the linear information transformation, and ReLU()($\cdot$)( ⋅ ) function transforms the graph information nonlinearly. $\bm{\gamma}$ is also a trainable parameter to measure the overall influence of related stock market indices. The GNNHAR model uses spatial GNNs to capture the nonlinear relationships between stock market indices and generates more accurate RV forecasts compared to the HAR model (Zhang et al., 2025). In addition, the GNNHAR model applies the same set of weights $\{\mathbf{W}^{(l)}\}_{l=0}^{L}$ to different parts of the volatility network to model the volatility spillover effect. This design ensures the GNNHAR model is scalable to large number of time series.
There are some other HAR model extensions such as the VHAR and HAR-KS models. These two models extend the linear combination of the HAR model to include historical volatility patterns of all entities in the system to generate RV forecasts for each target entity. However, they are less powerful than the GNNHAR model because: (a). they are less informative about the structure of the volatility network, and (b). they oversimplify the volatility spillover effect by assuming it is linear. Moreover, these models have potential scalability issue, as the number of parameters in both the VHAR and HAR-KS models grows quadratically as the number of relevant stock market indices increases. For further details, readers can refer to Bubák et al. (2011) for the VHAR model and Liang et al. (2020) for the HAR-KS model.
3 Methodology and Implementation
The proposed GSPHAR framework is introduced in this section. There are two proposed models, the basic GSPHAR model and the dynamic-GSPHAR (d-GSPHAR) model, under the GSPHAR framework. The d-GSPHAR model uses a simple yet effective method to model the dynamic volatility interconnections.
@@ 3.1 Model Formulation
In this section, the proposed GSPHAR model is introduced. The general workflow of the GSPHAR model is described in Figure 1 below:
@@
Now we present the detailed step-by-step development of the proposed framework. Suppose that the RV time series $\{\mathbf{v}_{t}\in\mathbb{R}^{N}\}_{t=1}^{T}$ consist of the RV time series of $N$ different stock market indices from time $t=1$ to time $t=T$. In the first step, the volatility spillover graph should be determined. Methods such as the Pearson correlation method and the Graphical LASSO (GLASSO) method (Friedman et al., 2007) only produce symmetric matrices to represent pairwise connections between stock markets. These methods do not capture the directional transmission of volatility. Influential stock markets tend to transmit more volatility to impressionable stock markets, whereas impressionable stock markets have minimal influence on influential stock markets. As a result, an alternative approach, the Diebold and Yilmaz (DY) framework (Diebold and Yilmaz, 2012), is applied to account for these asymmetric and directional effects. Specifically, the net pairwise volatility spillover effect matrix is calculated under the DY framework and is used as the adjacency matrix $\mathbf{A}^{\text{DY}}$ of the volatility spillover graph. $\mathbf{A}^{\text{DY}}$ can effectively highlight the dominant channels of volatility transmission, which allows researchers to pinpoint the markets that exert the most significant influence over others. The detailed calculation process of the net pairwise volatility spillover effect matrix can be found in Appendix B.
@@
Based on the adjacency matrix $\mathbf{A}^{\text{DY}}$, the normalized magnet Laplacian matrix of the volatility spillover graph can be calculated based on the process described from Equation (2) to Equation (7) with $\mathbf{A}=\mathbf{A}^{\text{DY}}$. To learn from the directional volatility spillover effect, the hyperparameter $q$ of the magnetic Laplacian should be positive. Since there is only one directed edge exists between any two connected nodes in the net pairwise volatility spillover graph, $q=0.25$ is chosen for better clarity and interpretability. Since $q=0.25$, the eigendecomposition of the magnetic Laplacian $\widetilde{\mathbf{L}}_{m}^{(0.25)}$ returns two eigenvector matrices, $\mathbf{U}_{m}$ and $\mathbf{U}_{m}^{\dagger}$, in the complex space and one eigenvalue matrix, $\bm{\Lambda}_{m}$, containing non-negative real numbers in its diagonal as shown in Equation (8). Thus, the RV time series input $[\mathbf{v}_{t-1},\mathbf{v}_{t-2},\ldots,\mathbf{v}_{t-22}]\in\mathbb{R}^{N\times 22}$ can be projected to the spectral domain through the GFT:
@@
Here, each projected value in $\{\widetilde{\mathbf{v}}_{t-p}\}_{p=1}^{22}$ is complex valued and can be written as: $\widetilde{\mathbf{v}}_{t-p}=\widetilde{\mathbf{v}}^{\text{r}}_{t-p}+i\widetilde{\mathbf{v}}^{\text{i}}_{t-p}$. The projected RV data merges the volatility information between different stock markets under the guidance of the volatility spillover graph. The GSPHAR model is designed to consistently process information from both the real and imaginary domains, following a similar approach as the HAR model.
The HAR family of models always applies an equal-weight average to construct the volatility pattern features based on the historical RV data. This assumption can be potentially improved. For example, under the simple design of the conventional HAR model in Equation (11), $\text{RV}_{t-2}$ and $\text{RV}_{t-5}$ always exert the same influence on the forecast $\text{RV}_{t}$. This equal weighting may not be able to sufficiently account for the dynamics of financial markets where the influence of past volatility on future volatility typically diminishes over time. This diminishing influence not only exists between short-term, mid-term and long-term, but also possibly occurs within different periods. To handle such potential insufficiency, a set of graph Fourier basis-specific convolution filters with learnable convex weights is applied to aggregate past information in the mid-term and long-term:
@@
Hereafter, the HAR model is built for the real and imaginary parts of the spectral RV inputs, respectively. For clarity, let $\widetilde{\mathbf{V}}_{t-22:t-1}=[\mathbf{1},\widetilde{\mathbf{v}}_{t-1},\widetilde{\mathbf{v}}_{t-5:t-1},\widetilde{\mathbf{v}}_{t-22:t-1}]\in\mathbb{R}^{N\times 4}$ to denote a collection of intercept terms and past short-term, mid-term and long-term spectral volatility patterns. The matrix $\mathbf{W}\in\mathbb{R}^{4\times 1}$ contains the linear transformation parameters used in the HAR model. Here, superscript ‘r’ is used to denote the real component and superscript ‘i’ represents the imaginary component. The same set of HAR coefficients is applied to both real and imaginary parts to aggregate the historical spectral volatility pattern. The real and imaginary parts of the spectral RV forecast $\widetilde{\mathbf{v}}_{t}$ are formulated as:
$\begin{split}\widetilde{\mathbf{v}}^{\text{r}}_{t}&\;=\widetilde{\mathbf{V}}_{t-22:t-1}^{\text{r}}\mathbf{W},\\ \widetilde{\mathbf{v}}^{\text{i}}_{t}&\;=\widetilde{\mathbf{V}}_{t-22:t-1}^{\text{i}}\mathbf{W}.\end{split}$ (15)
Set $\widetilde{\mathbf{v}}_{t}=\widetilde{\mathbf{v}}^{\text{r}}_{t}+i\widetilde{\mathbf{v}}^{\text{i}}_{t}$, which is still complex. $\widetilde{\mathbf{v}}_{t}$ is projected back to the spatial domain through the IGFT:
@@
The complex valued $\overline{\mathbf{v}}_{t}$ is divided into its real and imaginary parts again: $\overline{\mathbf{v}}_{t}=\overline{\mathbf{v}}^{\text{r}}_{t}+i\overline{\mathbf{v}}^{\text{i}}_{t}$. In the spatial domain, the information of the real part and the imaginary part of $\overline{\mathbf{v}}_{t}$ is aggregated by a $3$ layer neural network $NN(\cdot)$ to output the final forecast: $\widehat{\mathbf{v}}_{t}=NN(\overline{\mathbf{v}}^{\text{r}}_{t},\overline{\mathbf{v}}^{\text{i}}_{t})$. The shallow neural network can merge the volatility information in the real and imaginary domains in a nonlinear way, which effectively enhances the expressiveness of the proposed GSPHAR model. Furthermore, the growth rate of the parameters remains linear with respect to the number of stock market indices included. This ensures that the GSPHAR model is not only powerful, but also scalable to large volatility systems.